OpenStudy (anonymous):

Show that the function is a solution of the differential equation y = xlnx - 4x Differential equation: x + y - xy' = 0

OpenStudy (anonymous):

compute the derivative of \(x\ln(x)-4x\), replace \(y'\) by that derivative, and you will see that you get 0

OpenStudy (kirbykirby):

Find the derivative of your function first (to find y'): y' = (1)lnx + x(1/x) - 4 = lnx+1-4 = lnx - 3 Now you substitute it into the differential equation: x + (xlnx - 4x) -x(lnx - 3) = x + xlnx - 4x -xlnx + 3x = x - x = 0

OpenStudy (anonymous):

\[y'=1+\ln(x)-4\] now \[x+y-x(1+\ln(x)-4)=0\] is true

OpenStudy (anonymous):

Thank you so much guys! My teacher assigned tons of problems this morning and I have been staring at the stuff all day so this is really helping me out. :)